OpenStudy (anonymous):

Find the number c that satisfies the conclusion of the Mean Value Theorem. f(x)= x/x+2 [1,4]

OpenStudy (anonymous):

Okay, so the MVT states that at some point c in this interval, f'(c) = [f(b) - f(a)]/(b-a). Set a = 1 and b = 4. You get f(a) = f(1) = (1/3) f(b) = f(4) = (2/3) Plugging these in, you would get f'(c) = (1/9). f'(c) would equal 2/(c+2)^2. Then solve the quadratic equation.

OpenStudy (anonymous):

Anyone confirm?

OpenStudy (anonymous):

how would i solve it?

OpenStudy (anonymous):

Use the quadratic formula and the answer should be about 2.243.

OpenStudy (anonymous):

oh, thank you!!!!!!!!!

OpenStudy (zarkon):

\[\frac{1}{9}=\frac{2}{(c+2)^2}\] \[9=\frac{(c+2)^2}{2}\] \[\sqrt{9\cdot2}=c+2\] \[c=3\sqrt{2}-2\]

OpenStudy (anonymous):

thanks!